OAB company intends to distribute dividends to its stockholders amounting to BD12 per share next year....

  1. OAB company intends to distribute dividends to its stockholders amounting to BD12 per share next year. The amount represents 15% return of investor’s money. However, the company as part of its strategy, plow 40% of its earnings back to the company’s return on equity of 25%. Determine the value of the stock before plow-back and the value after plow-back.
Answer #1

Given D1=12

a. If there is no plow-back, retention rate=0. growth rate formula=retention rate*return on equity=0*25%=0

Value of the stock before plow-back=D1/(return-growth rate)=12/(15%-0)=80

b. if there is plow back,retention rate=40%. growth arte=40%*25%=10%

Value of the stock after plow-back=D1/(return-growth rate)=12/(15%-10%)=240